Two plano convex lens each of radius 100 are placed with their curved surfaces in contact with each other. Newton's rings are formed by a light of wavelength 6000A°. Find the distance between 10th and 20th fringe.
1
Expert's answer
2019-08-20T10:08:28-0400

We can write for the radius of dark ring

"r_{k}=\sqrt{kRλ} (1)"

In our two cases

"r_{10}=\sqrt{10Rλ} (2)"

"r_{20}=\sqrt{20Rλ} (3)"

The distance between 10th and 20th fringe is equal to

"Δr=\sqrt{20Rλ}-\sqrt{10Rλ} (4)"

Using (4) we get

Δr=8.83×10-3 m


Answer

8.83×10-3 m